OpenStudy (anonymous):

The length of a rectangular floor is 3 meters less than three times its width. If a diagonal of the rectangle is 13 meters, find the length and width of the floor. width m

OpenStudy (anonymous):

if length of rectangular floor is l, width is m and diagonal is n then l=3m-3 by Pythagoras theorem, l^2+m^2=n^2 (3m-3)^2+m^2=13^2 9m^2+9-18m+m^2=169 10m^2-18m-160=0 by solving, m=5 width-m=5 meters length=3m-3=12 meters